# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v3

@@ Abstract
[inst2]organization=Department of Radiology, Tokai University School of Medicine,addressline=143 Shimokasuya, city=Isehara, postcode=259-1193, state=Kanagawa, country=Japan
@@
We propose an unprecedented gastric cancer diagnostic support system for gastric X-rays, which can be taken by technicians and allow more people to be screened, rather than endoscopy, which is currently the mainstream but can only be performed by physicians. The system is based on a general deep learning-based object detection model and includes two novel technical proposals; a refined stochastic gastric image augmentation (R-sGAIA) and a hard boundary box learning (HBBT).
A R-sGAIA is a probabilistic gastric fold region enhancement method to provide more learning patterns for cancer detection models.
A HBBT is an efficient and versatile training method for common object detection models that allows the use of unannotated negative (i.e., healthy control) samples that cannot be used for training in conventional detection models. This reduces false positives and thus improves overall performance.
The sensitivity of the proposed system for gastric cancer (90.2%) exceeds that of the expert (85.5%), while 42.5% of the detected candidate box regions show cancerous lesions, with a processing time of 0.51 seconds/image. The proposed system showed 5.9 points higher on the F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. In short, the system quickly and efficiently shows the radiologist where to look, greatly reducing the radiologist’s workload.
@@ 1 Introduction
Gastric cancer is the third most deadly cancer in the world with a poor prognosis, especially in the advanced stages (Rawla and Barsouk, 2019). However, early prognosis before metastasis can improve outcome; therefore, early detection and appropriate treatment are important. Gastric cancer is primarily diagnosed by endoscopy or radiography. Endoscopy reported a sensitivity (SE) of 95.4%, which is superior to other methods such as gastric radiography or X-ray at 85.5% (Hamashima et al., 2013). However, endoscopy can only be performed by a physician, and the time and cost involved limits the number of people who can be screened. Gastric X-ray, on the other hand, can be performed by radiographers and provides a large number of images, making it suitable for screening. This is a major advantage over endoscopic diagnosis. In Japan, a mass screening system was established, and in 2019, at least 3.87 million people underwent gastric X-ray for gastric cancer screening, with 183,000 people indicated as needing further testing, and 2,553 cases of gastric cancer were ultimately found. On the other hand, only 374,000 people underwent endoscopy for the same purpose and were confirmed (Matsuura, 2023). Although the effectiveness of endoscopy has led to a shift to this diagnostic method, the development of superior diagnostic support technology for gastric radiography would significantly contribute to the early detection and reduction of mortality from gastric cancer.
@@
The gastric X-ray images obtained are later read by the doctor to diagnose the shape of the stomach and the atrophy of the mucosa membrane (Uemura et al., 2001; Roder, 2002; Correa et al., 2004). Another advantage of X-rays is that they can show minute irregularities that are indicative of stage IIc lesions (early stage tumor with a slight indentation) – slightly elevated component within the indentation that lies beneath the indentation. However, this is an area that requires a high level of reading skill, an area that many gastrointestinal angiographic studies are unable to diagnose. In other words, when performed by a skilled physician, the possibility of being able to assess the extent of cancerous invasion deep within the surface of an early stage cancer based on changes in the microscopic structure of the surface is high. Hence, the limitations of this diagnostic method are that it requires extensive physician experience in diagnosis and is particularly difficult to detect early-stage cancers, resulting in a cancer detection capacity that is inferior to that of endoscopy (Hibino et al., 2023). Therefore, there is a need to realize an automatic screening assistance system that can accurately and quickly detect gastric abnormalities from gastric X-ray images in clinical settings.
@@
Healthy gastric folds are thin, smooth, and parallel; however, their thickness and size can be altered by gastric diseases, such as inflammation, infection, ulcers, and tumors, resulting in irregular bends, depressions, and tears in the lining mucosa (Correa et al., 2004). Most of the research on diagnostic support using gastric X-rays aimed at detecting gastritis, the stage before gastric cancer or the presence of Helicobacter pylori infection, which is believed to be the cause of gastric cancer (Cooper et al., 1990; Kita, 1992; Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015, 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Among these proposals, methods based on deep learning represented by convolutional neural networks (CNNs) have reported excellent results in the detection of gastritis. The main advantage of these deep learning techniques is that they can automatically capture efficient features for the target task, which is promising in this area where disease features are difficult to define. However, studies directly detecting gastric cancer from radiographs are limited despite of the significance of gastric cancer diagnosis in terms of life expectancy. This may be because gastric cancer is more difficult to diagnose, especially in the early stages, compared to gastritis, where the lesion often expands and develops to some degree. Reportedly, it is difficult to distinguish gastritis from gastric cancer even when using the latest machine learning techniques on endoscopic images where signs of the disease are easily visible (Horiuchi et al., 2020).
@@
In the field of computer vision research, CNN-based object detection methods (e.g., Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon and Farhadi, 2018), EfficientDet (Tan et al., 2020), and YOLOv7 (Wang et al., 2023)), which can simultaneously detect and classify the location of multiple objects in images, have been proposed and widely used. Because these object detection methods can identify the regions involved, they have high explanatory power for the predicted results, as opposed to simply determining benign or malignant on an image-by-image basis. These features are extremely important in medical applications (Hirasawa et al., 2018; Li et al., 2018; Okamoto et al., 2019; Laddha et al., 2019; Zhang et al., 2019). In a study of gastric X-ray images, Laddha et al. (2019) used YOLOv3 to automatically detect gastric polyps, achieving a mean average precision of 0.916. On the other hand, these object detection methods require annotation of lesion locations prior to training, which is extremely expensive. The aforementioned gastritis detection studies (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020) used a patch-based CNN that only needed to be annotated on an image-by-image basis to avoid high annotation costs while fully understanding the explanatory advantages provided by the object detection model. However, in gastric cancer, where diagnostic difficulties and outcomes are more severe, it is critical that physicians are able to interpret the results of diagnostic support systems.
With this motivation, we proposed a diagnostic support system that directly targets gastric cancer using X-ray images, which have rarely been studied, and provides high explanatory power for the results (Okamoto et al., 2019). Thanks to the proposed stochastic gastric image augmentation (sGAIA) based on medical knowledge, this system was able to represent candidate regions of gastric cancer in a boundary box with high accuracy. The recall (= SE) and precision of the detected bounding boxes were 92.3% and 32.4%, respectively, in an evaluation using images of patients different from the training set. This is about 7% higher recall than that achieved by doctors (85.5%), and the box-by-box evaluation confirmed that the ratio of true to false detections was 1:3, which is sufficient for practical use. The object detection model trained with the sGAIA has achieved a practical level of accuracy; however, further performance improvements are desirable.
Currently, widely used object detection methods have the disadvantage of not being able to explicitly use negative (i.e., healthy control) data for learning, in addition to the high cost of annotation. When such models are used for lesion detection, only images with lesions are used for training; control images without lesions cannot be used. There is much room for improvement here, which is one of the main themes of the work in this study. When dealing with medical images, data are expensive, and we believe that actively using healthy control images that do not contain lesions for training can improve the overall ability of the model to detect lesions. This is equivalent to reducing the cost of annotation to achieve the same performance.
@@
In this study, we proposed a practical gastric cancer screening system. The system includes two new proposals: refined stochastic gastric image augmentation (R-sGAIA), which is an improvement of sGAIA (Okamoto et al., 2019), and hard boundary box training (HBBT, enables the learning of negative or healthy control images), which have not been previously explicitly utilized for training in object detection neural networks. To suppress the high-confidence false positives predicted by the object detector, HBBT registers each of these boxes as a “hard-sample” class and retrains the model until the predetermined convergence conditions are met. We investigated 4,608 gastric X-ray images obtained from 140 patients in a clinical setting and confirmed the practicality and effectiveness of the proposed system by evaluating gastric cancer detection performance with a patient-based cross-validation strategy.
@@
    The SE of the proposed system for gastric cancer (90.2%) exceeds that of the expert (85.5%), while 42.5% of the detected candidate box regions show cancerous lesions, with a processing time of 0.51 s/image. The proposed system showed 5.9 points higher on the F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. In short, the system quickly and efficiently shows the radiologist where to look, greatly reducing the radiologist’s workload.
@@ 2.1 Diagnostic support using gastric X-ray images
In the 1990s, methods for effectively extracting the pattern of gastric folds by applying binarization as a preprocessing step and for estimating the gastric region based on the position of the barium reservoir were proposed (Cooper et al., 1990; Kita, 1992). In the 2010s, more empirical studies focused on gastric folds began to emerge (Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015). Ishihara et al. (2015) calculated numerous statistical features (7,760 per case) for H. pylori infection, a major cause of gastric cancer, by focusing on post-infection mucosal patterns and gastric folds. Their support vector machine based on the data of 2,100 patients, 8 per person, yielded an SE of 89.5% and a specificity (SP) of 89.6%.
Recently, taking advantage of deep learning techniques, research has been conducted on the detection of gastritis using patch-based CNNs on gastric X-ray images (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Although these methods targeted gastritis, which is easier to diagnose than gastric cancer, they achieved high detection accuracy and were considered practical in diagnostic assistance. Kanai et al. (2020) used a patch-based CNN to detect H. pylori-associated chronic atrophic gastritis. They trained image patches of the inside and outside of the stomach and made a diagnosis by a majority vote of each estimated result in the stomach region. Manual annotation of regions of interest (ROIs) prior to training and self-learning to increase the number of training patches ultimately yielded a harmonic mean SE and SP of 95.5% for the test data of different patients.
While excellent screening results have been achieved for gastritis, the detection of more serious and difficult to diagnose gastric cancers has not been well studied. In the case of gastric cancer, it is also important to provide an explanation for the results. The proposed method is a significant improvement of our previous work, a gastric cancer detection system that can present explainable results (Okamoto et al., 2019), with two new technical proposals.
@@ 2.2 Diagnostic support research for endoscopic images
In recent years, many methods for analyzing narrow-band endoscopy images to assist in the diagnosis of gastric cancer using deep learning technology have been proposed (Kanayama et al., 2019; Wu et al., 2019; Horiuchi et al., 2020; Li et al., 2020; Gong et al., 2023).
@@
Wu et al.’s (2019) classic VGG16 or ResNet-50 models achieved an SE of 94.0% and an SP of 91.0% based on 24,549 images for diagnosing early gastric cancer. Li et al.’s (2020) Inception-v3 model achieved an SE of 91.2% and SP of 90.6% based on 2,429 images of gastric magnification endoscopy with narrow-band imaging. Horiuchi et al. (2020) trained GoogLeNet with 1,499 images of gastric cancer and 1,078 images of gastritis and achieved an SE of 95.4%, an SP of 71.0%, and an area under the receiver operating characteristic (ROC) curve (AUC) of 0.852. Gong et al. (2023) developed a Convolution and Relative Self-Attention Parallel Network based on 3,576 endoscopic images taken from 205 patients. They reported achieving F1 = 0.948 in their experiments with a 7:3 split of training and test images. Although these methods produce excellent results, none of them have statements that clearly separate the training dataset from the patients for whom the images in the evaluation set were taken. In other words, it is possible that the same patient images are mixed in the training and test data. In this case, the test data, which were very close to the training data, were evaluated, resulting in a score that was much higher than the original ability.
@@
Kanayama et al. (2019) constructed a diagnostic support system using 129,692 gastrointestinal endoscopy images with sophisticated CNN-based networks, including one synthesizer network and two different types of discriminator networks .Furthermore, the images of the same patient were not divided into training and evaluation sets; that is, the performance were evaluated appropriately. They reported an average precision (AP) of 0.596 in cancer detection performance. Although the numerical performance of this method appears inferior, it is based on a newer and more sophisticated ML model than previously reported methods and is considered to represent a true diagnostic capability for this problem based on current machine learning techniques. They generated 20,000 high-resolution lesion images from 1,315 narrow-field-of-view images of the lesion using their own generative adversarial network (Goodfellow et al., 2014) and added these images to the training set. As a result, the AP reported an improvement in cancer detection performance to 0.632. Shibata et al. (2020) constructed and rigorously evaluated a diagnostic model using Mask R-CNN on 1,208 and 533 images each obtained from 42 healthy subjects and 93 unhealthy subjects (including gastric cancer), respectively. Their model achieved an SE of 96% for all tumor types, a false positive of 0.1/image, and F1 = 0.71 for gastric cancer. Ahmad et al. (2023) trained and evaluated a larger dataset than previous studies (27,200 training and 6,800 testing) with a unique improved method that added Squeeze and Excitation attention blocks to YOLO-v7 (Wang et al., 2023), achieving F1 = 0.71 (precision = 72%, recall = 69%). While we could not confirm if proper data partitioning was done, it is important to note that the number of images used for training and evaluation was sufficiently large compared to other studies. In addition, the study used more sophisticated methods, with results different from the previous studies (about 90% in F1 and higher) where substantial data leakage was considered to have occurred. Lastly, the results were comparable to other reports (Shibata et al., 2020) where the distinction was clearly split. It can be inferred that the data division was appropriate.

### v4

@@ Abstract
[inst2]organization=AI Development Department, Aillis Inc., addressline=2-2-1 Yaesu, city=Chuo, postcode=104-0028, state=Tokyo, country=Japan
[…]
[inst3]organization=Department of Radiology, Tokai University School of Medicine,addressline=143 Shimokasuya, city=Isehara, postcode=259-1193, state=Kanagawa, country=Japan
@@
We propose an unprecedented gastric cancer diagnostic support system for gastric X-rays, which can be performed by technicians, allowing more people to be screened compared to endoscopy, the current mainstream method that can only be performed by physicians. The system is based on a general deep learning-based object detection model and introduces two novel techniques: refined stochastic gastric image augmentation (R-sGAIA) and hard boundary box training (HBBT).
A R-sGAIA is a probabilistic gastric fold region enhancement method that provides more learning patterns for cancer detection models.
A HBBT is an efficient and versatile training method for common object detection models, enabling the use of unannotated negative (i.e., healthy control) samples that cannot be used in conventional detection models. This reduces false positives and improves overall performance.
The proposed system achieves a sensitivity of 90.2% for gastric cancer, exceeding that of an expert (85.5%), with 42.5% of the detected candidate box regions showing cancerous lesions and a processing time of 0.51 seconds per image. It also achieves a 5.9-point higher F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. The system efficiently guides radiologists on where to focus, significantly reducing their workload.
Our R-sGAIA and HBBT are abailable at https://github.com/IyatomiLab/RsGAIA_HBBT/.
@@ 1 Introduction
Gastric cancer is the third most deadly cancer globally, with a poor prognosis, especially in advanced stages (Rawla and Barsouk, 2019). However, early detection before metastasis can significantly improve outcomes, making early diagnosis and appropriate treatment crucial. Gastric cancer is primarily diagnosed through endoscopy or radiography. Endoscopy reports a sensitivity (SE) of 95.4%, which is superior to other methods, such as gastric radiography or X-ray, at 85.5% (Hamashima et al., 2013). However, endoscopy can only be performed by a physician, and the time and cost involved limit the number of people who can be screened. In contrast, gastric X-rays can be performed by radiographers, generating a large number of images, which makes this method suitable for mass screening – a significant advantage over endoscopic diagnosis. In Japan, a mass screening system was established, and in 2019, at least 3.87 million people underwent gastric X-rays for gastric cancer screening. Of these, 183,000 people were indicated for further testing, and 2,553 cases of gastric cancer were ultimately detected. By comparison, only 374,000 people underwent endoscopy for the same purpose (Matsuura, 2023). Although endoscopy’s effectiveness has led to a shift toward this diagnostic method, developing superior diagnostic support technology for gastric radiography would significantly contribute to the early detection and reduction of mortality from gastric cancer.
@@
Gastric X-ray images are later interpreted by doctors to diagnose the stomach’s shape and mucosal atrophy (Uemura et al., 2001; Roder, 2002; Correa et al., 2004). An additional advantage of X-rays is their ability to reveal minute irregularities indicative of stage IIc lesions (early-stage tumors with slight indentations) – a subtle elevation within the indentation that lies beneath it. However, accurately reading these images requires a high level of skill, which many gastrointestinal radiologists may lack. When performed by a skilled physician, there is a high likelihood of assessing the extent of cancerous invasion deep within the surface of an early-stage cancer based on changes in the microscopic structure of the surface. The limitations of this diagnostic method lie in its dependence on extensive physician experience and the particular difficulty in detecting early-stage cancers, leading to a lower cancer detection capacity than that of endoscopy (Hibino et al., 2023). Therefore, there is a pressing need for an automated screening assistance system that can accurately and promptly detect gastric abnormalities from gastric X-ray images in clinical settings.
@@
Healthy gastric folds are thin, smooth, and parallel; however, their thickness and size can be altered by gastric diseases such as inflammation, infection, ulcers, and tumors, leading to irregular bends, depressions, and tears in the lining mucosa (Correa et al., 2004). Most research on diagnostic support using gastric X-rays has focused on detecting gastritis, the precursor to gastric cancer, or the presence of Helicobacter pylori infection, which is believed to be a major cause of gastric cancer (Cooper et al., 1990; Kita, 1992; Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015, 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Among these studies, methods based on deep learning, particularly convolutional neural networks (CNNs), have shown excellent results in detecting gastritis. The main advantage of these deep learning techniques is their ability to automatically capture efficient features for the target task, which is promising in areas where disease features are difficult to define. However, studies directly detecting gastric cancer from radiographs are limited, despite the significance of gastric cancer diagnosis in terms of life expectancy. This may be because gastric cancer, especially in its early stages, is more difficult to diagnose compared to gastritis, where lesions often expand and develop to some degree. Reportedly, even with the latest machine learning techniques, distinguishing between gastritis and gastric cancer on endoscopic images—where signs of the disease are more visible – remains challenging (Horiuchi et al., 2020).
@@
In the field of computer vision, CNN-based object detection methods (e.g., Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon and Farhadi, 2018), EfficientDet (Tan et al., 2020), and YOLOv7 (Wang et al., 2023)), which can simultaneously detect and classify the location of multiple objects in images, have been proposed and widely used. These methods are particularly valuable in medical applications because they not only detect regions of interest but also provide high explanatory power for the predicted results, as opposed to simply classifying images as benign or malignant (Hirasawa et al., 2018; Li et al., 2018; Okamoto et al., 2019; Laddha et al., 2019; Zhang et al., 2019). For instance, in a study on gastric X-ray images, Laddha et al. (2019) used YOLOv3 to automatically detect gastric polyps, achieving a mean average precision of 0.916. However, these object detection methods require precise annotation of lesion locations prior to training, which is extremely costly. To avoid high annotation costs while still benefiting from the explanatory advantages of object detection models, previous studies on gastritis detection (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020) used patch-based CNNs, which only required image-level annotations. In the context of gastric cancer, where diagnostic challenges and outcomes are more severe, it is critical that physicians are able to interpret the results provided by diagnostic support systems.
@@
With this motivation, we proposed a diagnostic support system that directly targets gastric cancer using X-ray images, which have rarely been studied, and offers high explanatory power for the results (Okamoto et al., 2019). Thanks to the proposed stochastic gastric image augmentation (sGAIA) based on medical knowledge, this system accurately represented candidate regions of gastric cancer within bounding boxes. In an evaluation using images from patients not included in the training set, the recall (SE) and precision of the detected bounding boxes were 92.3% and 32.4%, respectively. This recall is about 7% higher than that achieved by doctors (85.5%), and a box-by-box evaluation confirmed a true-to-false detection ratio of 1:3, which is sufficient for practical use. While the object detection model trained with sGAIA has reached a practical level of accuracy, further performance improvements are desirable. Currently, widely used object detection methods have the disadvantage of not being able to explicitly use negative (i.e., healthy control) data for training, in addition to the high cost of annotation. When such models are used for lesion detection, only images with lesions are used for training, while control images without lesions are excluded. This limitation presents an opportunity for improvement, which is a key focus of this study. In medical imaging, data are expensive, and we believe that actively using healthy control images for training can enhance the model’s overall ability to detect lesions, effectively reducing annotation costs while maintaining performance. In this study, we proposed a practical gastric cancer screening system that includes two new techniques: refined stochastic gastric image augmentation (R-sGAIA), an improvement of sGAIA (Okamoto et al., 2019), and hard boundary box training (HBBT), which enables the use of negative or healthy control images for training, a feature not previously utilized in object detection neural networks. To suppress high-confidence false positives predicted by the object detector, HBBT registers each of these boxes as a “hard-sample” class and retrains the model until predetermined convergence conditions are met. We investigated 4,724 gastric X-ray images obtained from 145 patients in a clinical setting and confirmed the practicality and effectiveness of the proposed system by evaluating gastric cancer detection performance using a subject-based five-fold cross-validation strategy.
@@
    The SE of the proposed system for gastric cancer (90.2%) exceeds that of the expert (85.5%), while 42.5% of the detected candidate box regions show cancerous lesions, with a processing time of 0.51 s/image. It is 5.9 points higher on the F1 score compared to methods using the same object detection model and state-of-the-art data augmentation. In short, the system efficiently shows the radiologist where to look, greatly reducing the radiologist’s workload.
  • 5.
    Our R-sGAIA and HBBT are abailable at https://github.com/IyatomiLab/RsGAIA_HBBT/.
@@ 2.1 Diagnostic support using gastric X-ray images
In the 1990s, methods were proposed for effectively extracting the pattern of gastric folds by applying binarization as a preprocessing step and estimating the gastric region based on the position of the barium reservoir (Cooper et al., 1990; Kita, 1992). In the 2010s, more empirical studies began to focus on gastric folds (Minemoto et al., 2010; Nagano and Matsuo, 2010; Abe et al., 2014; Ishihara et al., 2015). Ishihara et al. (2015) calculated numerous statistical features (7,760 per case) for H. pylori infection, a major cause of gastric cancer, by analyzing post-infection mucosal patterns and gastric folds. Their support vector machine, trained on data from 2,100 patients (with eight images per person), achieved a SE of 89.5% and a SP of 89.6%.
More recently, deep learning techniques have been applied to the detection of gastritis using patch-based CNNs on gastric X-ray images (Ishihara et al., 2017; Togo et al., 2019; Kanai et al., 2019, 2020). Although these methods targeted gastritis, which is easier to diagnose than gastric cancer, they achieved high detection accuracy and were considered practical for diagnostic assistance. Kanai et al. (2020) used a patch-based CNN to detect H. pylori-associated chronic atrophic gastritis. They trained image patches from both the inside and outside of the stomach and made a diagnosis by taking a majority vote of each estimated result in the stomach region. Manual annotation of regions of interest (ROIs) prior to training, along with self-learning to increase the number of training patches, ultimately yielded a harmonic mean SE and SP of 95.5% for the test data from different patients. While excellent screening results have been achieved for gastritis, the detection of more serious and difficult-to-diagnose gastric cancers has not been well studied. In the case of gastric cancer, it is also crucial to provide an explanation for the results. The proposed method significantly improves upon our previous work on a gastric cancer detection system that can present explainable results (Okamoto et al., 2019), incorporating two new technical advancements.
@@ 2.2 Diagnostic support research for endoscopic images
In recent years, many methods for analyzing narrow-band endoscopy images to assist in diagnosing gastric cancer using deep learning technology have been proposed (Kanayama et al., 2019; Wu et al., 2019; Horiuchi et al., 2020; Li et al., 2020; Gong et al., 2023).
@@
Wu et al. (2019) used classic VGG16 and ResNet-50 models, achieving a SE of 94.0% and a SP of 91.0% based on 24,549 images for diagnosing early gastric cancer. Li et al. (2020) employed an Inception-v3 model, achieving an SE of 91.2% and an SP of 90.6% based on 2,429 images of gastric magnification endoscopy with narrow-band imaging. Horiuchi et al. (2020) trained a GoogLeNet model with 1,499 images of gastric cancer and 1,078 images of gastritis, achieving an SE of 95.4%, an SP of 71.0%, and an area under the receiver operating characteristic (ROC) curve (AUC) of 0.852. Gong et al. (2023) developed a Convolution and Relative Self-Attention Parallel Network using 3,576 endoscopic images from 205 patients, reporting an F1 score of 0.948 in their experiments with a 7:3 split of training and test images. While these methods produce excellent results, none of them explicitly state whether the training dataset is completely separate from the patients whose images are in the evaluation set. This raises the possibility that images from the same patients were included in both the training and test data, potentially leading to inflated performance scores due to the close similarity between the two datasets.
@@
In contrast, Kanayama et al. (2019) constructed a diagnostic support system using 129,692 gastrointestinal endoscopy images with sophisticated CNN-based networks, including one synthesizer network and two different types of discriminator networks. Importantly, they ensured that images from the same patient were not divided between the training and evaluation sets, thus evaluating performance appropriately. They reported an average precision (AP) of 0.596 for cancer detection. Although the numerical performance of this method may seem lower, it is based on a newer and more sophisticated machine learning model than those used in previously reported methods, representing a more accurate diagnostic capability. They further enhanced their model by generating 20,000 high-resolution lesion images from 1,315 narrow-field-of-view images using their own generative adversarial network (Goodfellow et al., 2014) and adding these to the training set, improving the AP in cancer detection to 0.632. Shibata et al. (2020) constructed and rigorously evaluated a diagnostic model using Mask R-CNN on 1,208 images from 42 healthy subjects and 533 images from 93 unhealthy subjects (including gastric cancer). Their model achieved an SE of 96% for all tumor types, a false positive rate of 0.1/image, and an F1 score of 0.71 for gastric cancer. Ahmad et al. (2023) trained and evaluated a larger dataset than in previous studies (27,200 training and 6,800 testing images) with a novel approach that added Squeeze and Excitation attention blocks to YOLOv7 (Wang et al., 2023), achieving an F1 score of 0.71 (precision = 72%, recall = 69%).
While it is unclear whether proper data partitioning was employed, it is important to note that the number of images used was significantly larger than in other studies. Additionally, the use of more sophisticated methods yielded results different from earlier studies (which reported F1 scores around 90% or higher) where substantial data leakage was suspected. Given that the results were comparable to other reports (Shibata et al., 2020) with clearly distinct data partitions, it can be inferred that the data division was likely appropriate.